OpenStudy (anonymous):

Solve algebraically for x: x plus 2 over 6 equals 3 over x minus 1

OpenStudy (anonymous):

\[\frac{x+2}{6}=\frac{3}{x-1}\] \[(x+2)(x-1)=3\times 6\] is a start